OpenStudy (anonymous):

I don' understand this question If MI = 15 m, SM = 20 m, and TM = 9 m, what is the value of RM? 12 m 18 m 20 m 30 m

OpenStudy (anonymous):

do u have a diagram??

OpenStudy (anonymous):

that is the picture it come with

OpenStudy (anonymous):

erm it use pythagoras for IT....do you know anything about the relationship between parallel hypotenuses ???

OpenStudy (anonymous):

no not really i am just really confused in how to find the answer

OpenStudy (anonymous):

i think it is 18 but idk if that is right

OpenStudy (reemii):

think of the triangle SRM as the little triangle ITM scaled by some factor. If you find the scale between two corresponding sides, then you can use that scale same for all other sides. MI corresponds to MS. the scale is 4/3. right?

OpenStudy (reemii):

(MI/3) * 4 = MS. then (MT/3)*4 = RM

OpenStudy (anonymous):

but the answers are 12 m 18 m- and i think this is the right one 20 m 30 m

OpenStudy (anonymous):

im think the answer is 12...based frm reemii explanation the length of IT= 3 root 34 if you multiply that by 20/15...u get SR once u have SR use pythagoras becos u have SM....to find MR

OpenStudy (anonymous):

ok i see what i did wrong thank you

OpenStudy (anonymous):

no probzz i actually learnt something myself!!

OpenStudy (reemii):

:D

OpenStudy (reemii):

he's right the answer is 12.

OpenStudy (anonymous):

shes right lol reemii im a girl LOL

OpenStudy (anonymous):

can you help me on one more

OpenStudy (anonymous):

What is the value of cos A?

OpenStudy (anonymous):

diagram??

OpenStudy (anonymous):

the answers are a=12/15 b=9/15 c=9/12 d=12/9

OpenStudy (anonymous):

i believe that the answer is b

OpenStudy (anonymous):

use the cosine rule....the answer is 53.13

OpenStudy (anonymous):

so then was i correct

OpenStudy (anonymous):

yep!!